# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v1

@@ III-C3 Loss Formulation
where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value. $n$ is the totoal number of pixels in the image. The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
@@ IV Experiments
TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result.
@@
$JRCFA$ $FA$ Heyex Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
@@
$JRCFAF$ $FAF$ Heyex Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
@@
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images. Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
@@ IV-A1 Image Translation Training
We prepared a multi-modality dataset with over 150k un-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, the original images are ultra-wide-filed 135-degree images, which are out of the scope of this work (will be dicussed in section V-A). Therefore, we crop random regions from the center area of the retina to get 30-40-degree conventional field-of-view images and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
@@ IV-A3 Evaluation
Seven diverse datasets are used for evaluation. For the CF modality, we use the public STARE [78], ChaseDB1 [79] and HRF [16] datasets. For MC modality, we use the public IOSTAR [80] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF dataset consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). More information on the 7 datasets can be found in Table I.
@@ IV-B1 Translation Network
The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task. We use a U-Net [28] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^2] More details on the network structures we use for translation and segmentation will be provided in our codes.. For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images. We start with an initial learning rate of $42e-$ 4 and start decaying the learning rate linearly to 0 from epoch 50 to 100. Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay. $\lambda_{c}=$10.0 and $\lambda_{i}=$0.5.
@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 4. The weight $\lambda_{tc}$ and $\lambda_{ts}$ we use are 0.05 and 0.0002, respectively. All the testing images are first pad to square and resized to 768$\times$768 to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [81].
@@ IV-C Comparison with Finetuned SOTA Methods
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent picks of datasets, dataset splits and inconsistent choices of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). The baseline methods include U-Net [28], CS2-Net [30], W-Net [39], FA-Net [40], GT-DLA-dsHff [31] and WS-DMF [38]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsampling them to 768$\times$768 for training and testing. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, since topological learning is an important part of our work and the topological accuracy is important in retinal vessel segmentation, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [62] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [63]. All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better.
@@
Table I, Fig. 8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense. However, for some modalities there still exist non-neglectable performance gap, especially for the FA. As shown by an example in the second row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
@@ IV-D2 Image Translation Method
For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as [58] uses because we find on larger training set StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images. In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design. In general, the final segmentation performance by StyleDiffusion is the poorest (Fig. 11). Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig. 12) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
@@
As shown in Fig. 11, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 12). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generate a smoother translated image.
IV-D3 Topological Segmentation Loss Function
In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM. Table II and Fig. 13 shows that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only using $\mathcal{L}_{ts}$, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
@@ V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation is imperfect and has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to more false negative pixels and the thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function. As a result, performance gap still exists between our method and finetuned SOTA methods. Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view. While there is an emerging type of 135 degree ultra-wide-field retinal images which our work has not included so far. In addition, there is another 10-15 degree OCTA retinal image which is also commonly used but is not only different in field-of-view, but also in the imaging principle. These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
V-B Conclusion
In this paper, we proposal a foundational universal vessel segmentation model for multi-modality retinal images. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new track in the research of retinal vessel segmentation.
V-C Future Work
Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to be more robust to images with worse imaging quality. We also aim for systematic study of how to allow the model to be more robust to the outlier patterns which the segmentation training set does not include, especially for severe pathological patterns on the retina (currently we find the topology-aware segmentation method helps addressing this issue, but since we do not have organized dataset for evaluating the problem, we did not report this topic in this paper).

### v2

@@ III-C3 Loss Formulation
where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value. $n$ is the total number of pixels in the image. The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
@@ IV Experiments
TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result. Heidel refers to Heidelberg Spectralis.
@@
$JRCFA$ $FA$ Heidel Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
@@
$JRCFAF$ $FAF$ Heidel Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
@@
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view). Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR. Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
@@ IV-A1 Image Translation Training
We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, the original images are ultra-wide-filed 135-degree images, which are out of the scope of this work (will be dicussed in section V-A). Therefore, we crop random regions from the center area of the retina to get 30-40-degree conventional field-of-view images and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
@@ IV-A3 Evaluation
Seven diverse datasets are used for evaluation. For the CF modality, we use the public STARE [78], ChaseDB1 [79] and HRF [16] datasets. For MC modality, we use the public IOSTAR [80] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). More information on the 7 datasets can be found in Table I. In addition to the seven vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [81]. We provide preliminary insights in this paper (Fig. 1, Fig. 8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.
@@ IV-B1 Translation Network
The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task. We use a U-Net [28] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^2] More details on the network structures we use for translation and segmentation will be provided in our codes.. For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images. We start with an initial learning rate of 2e-4 and start decaying the learning rate linearly to 0 from epoch 50 to 100. Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay. $\lambda_{c}=$10.0 and $\lambda_{i}=$0.5.
@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 4. The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively. All the testing images are first padded to square and resized to 768$\times$768 to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [82].
@@ IV-C Comparison with Finetuned SOTA Methods
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). The baseline methods include U-Net [28], CS2-Net [30], W-Net [39], FA-Net [40], GT-DLA-dsHff [31] and WS-DMF [38]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsampling them to 768$\times$768 for training and testing. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, since topological learning is an important part of our work and the topological accuracy is important in retinal vessel segmentation, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [62] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [63]. All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better. For all of them reported metrics, we compute them at the original image size and averaged over all test images.
@@
The results in Table I, Fig. 8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense. However, for some modalities there still exist non-neglectable performance gap, especially for the FA. As shown by an example in the second row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
@@ IV-D2 Image Translation Method
For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [58] because we find that on larger training set, StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images. In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design. In general, the final segmentation performance by StyleDiffusion is the poorest (Fig. 11). Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig. 12) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
@@
As shown in Fig. 11, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 12). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image.
@@
IV-D3 Topological Segmentation Loss Function
In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM. Table II and Fig. 13 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
@@ V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function. Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view. Our work does not include the ultra-wide-field 135 degrees retina images. Moreover, we does not consider the commonly-used 10-15 degree OCTA image, as its imaging principle is different. These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
V-B Conclusion and Future Work
In this paper, we propose a foundational universal vessel segmentation model for multi-modality retinal images. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new direction in the research of retinal vessel segmentation. Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to better predict the fine vanishing vessels.